OpenStudy (gygx72652):

HOW MANY INTEGERS FROM 4 TO 50, INCLUSIVE, ARE NEITHER MULIPLES OF 3 NOR 4?

OpenStudy (sshayer):

no. of multiples of 3 from 4 to 50 =15 no. of multiples of 4 from 4 to 50 =12 no. of multiples of 12 from 4 to 50 =4 no. of multiples of 3 or 4=15+12-4=23 total nos. from 4 to 50=50-3=47 no. of nos. which are not multiples of 3 and 4=47-23=24